3)Smiley Corporation's current sales and partial balance sheet are shown below.

This year
Sales $ 10,000
Balance Sheet: Liabilities
Accounts payable $ 1,500
Notes payable $ 2,500
Accruals $ 1,400
    Total current liabilities $ 5,400
Long-term bonds $ 2,000
    Total liabilities $ 7,400
Common stock $ 1,500
Retained earnings $ 2,000
    Total common equity $ 3,500
  Total liabilities & equity $ 10,900

Sales are expected to grow by 12% next year. Assuming no change in operations from this year to next year, what are the projected spontaneous liabilities? Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

Spontaneous liabilities are the liabilities which will increase without action between creditor and firm. For example, accounts payable, accrued expenses, accrued wages, accrued taxes.

However, notes payable is not a spontaneous liability as it needs an action between creditor and firm to raise this liability.

Projected sales = 10,000+(10,000*12%)

Projected sales = 11,200

Spontaneous liabilities = accounts payable + accruals

Spontaneous liabilities = 1,500+1,400 = $2,900

Projected spontaneous liabilities = spontaneous liabilities/sales*Projected sales

Projected spontaneous liabilities = 2,900/10,000*11,200

Projected spontaneous liabilities = $3,248